# Deployability-Centric Infrastructure-as-Code Generation: An LLM-based Iterative Framework

### v1

@@ VIII Related Work
Feedback has been widely adopted to enhance LLM outputs. For instance, Chen et al. [42] used natural language feedback to refine code completions, Huang et al. [44] introduced feedback signals during model training, and Yang et al [43] leveraged feedback for iterative output correction. Despite these advances, the use of feedback in IaC generation remains underexplored, particularly when considering the need for executable and deployable outputs. Our work differs by leveraging deployment-driven feedback to guide LLMs toward producing functionally valid IaC scripts.
IX Conclusion
@@
IaC generation holds significant promise for automating cloud infrastructure provisioning. We present IaCGen, a novel LLM-based deployability-centric framework that uses iterative feedback mechanism to generate IaC templates, and introduce DPIaC-Eval, a deployability-centric IaC template benchmark consists of 153 real-world scenarios that can evaluate syntax, deployment, user intent, and security. Our evaluation reveals that state-of-the-art LLMs initially performed poorly, with Claude-3.5 and Claude-3.7 achieving only 30.2% and 26.8% deployment success on the first attempt respectively. However, IaCGen transforms this performance dramatically: all evaluated models reach $>$90% passItr@25, with Claude-3.5 and Claude-3.7 achieving 98% success rate. Our work provides the first comprehensive assessment of deployability-centric IaC template generation and establishes a foundation for future research.

### v2

@@ 8.1. Feedback Mechanism of LLMs
Feedback mechanism has been widely adopted to enhance LLM outputs in SE tasks. For instance, Chen et al. (chen2023teaching) used natural language feedback to refine code completions, Huang et al. (huang2022large) introduced feedback signals during model training, and Yang et al. (yang2022generating) leveraged feedback for iterative output correction. Despite these advances, the use of feedback in IaC generation remains underexplored, particularly when considering the need for executable and deployable outputs.
Several studies have incorporated iterative feedback into LLM-based code generation (zhang2023repocoder; bi2024iterative; bouzenia2025you). However, these approaches typically provide only immediate error messages and problematic code snippets. Due to IaC’s declarative nature and complex resource interdependencies, addressing one error often introduces new issues elsewhere, as fixes frequently require coordinated changes across multiple template sections. Thus, in this study, we provide the complete conversation history to the LLM, including all previous errors and attempted solutions, enabling in-context learning (dong2022survey) and a more informed decision-making that avoids repeating previous mistakes.
8.2. LLMs for IaC Generation
Despite the growing importance of IaC in modern cloud development, research on LLMs’ capability to generate reliable infrastructure code remains limited. Kon et al. (kon2024iac) introduced the first benchmark for evaluating LLM-generated IaC, focusing primarily on syntactic validity and intent compliance across Terraform templates. They found LLMs achieved less than 20% success rates, which is substantially lower than general code generation tasks. Ragothaman (ragothaman2024optimizing) attempted to address deployability concerns but provided limited technical details about their validation process and relied heavily on manual testing procedures. While they identified the potential value of incorporating error feedback from failed templates into prompts or retrieval-augmented generation systems, they did not implement or evaluate such mechanisms. Similarly, Palavalli et al. (palavalli2024using) examined feedback loops for IaC generation, discovering that feedback effectiveness diminishes significantly after five iterations. However, their evaluation remained limited to syntax validation, and their feedback strategy only provided incorrect IaC template and basic error information.
Compared to existing research, our work advances this research area through three key contributions. We introduce the first comprehensive deployability-centric benchmark that validates templates in live cloud environments, capturing the complete spectrum of real-world deployment failures. Additionally, our framework substantially improves performance through deployment-driven iterative refinement. Finally, by prioritizing deployability, our approach provides the first practical evaluation of LLM capabilities on IaC generation in DevOps.
9. Conclusion
@@
IaC generation holds significant promise for automating cloud infrastructure provisioning in DevOps. We construct DPIaC-Eval, a deployability-centric IaC template benchmark consisting of 153 real-world scenarios; and propose IaCGen, a novel LLM-based deployability-centric framework that uses an iterative feedback mechanism to generate IaC templates. Our evaluation reveals that state-of-the-art LLMs initially performed poorly, with Claude-3.5 and Claude-3.7 achieving only 30.2% and 26.8% deployment success on the first attempt, respectively. However, IaCGen transforms this performance dramatically: all evaluated models reach >>90% passItr@25, with Claude-3.5 and Claude-3.7 achieving 98% success rate. This work provides the first comprehensive assessment of deployability-centric IaC template generation and establishes a foundation for future research.
Data Availability. The replication package contains two folders. The Data folder contains the benchmarks, and the Code folder contains the code of our IaCGen framework. Detailed descriptions of files can be found in the README.md file within each folder. The replication package is available at https://github.com/Tianyi2/IaCGen.